A container contains 2.5m 3 of Helium gas at 15 0 C and 98 Nm -2 pressures. What volume would this gas occupy at S.T.P? 


1
Expert's answer
2022-01-11T11:53:04-0500

The standard temperature is 273 K (0° Celsius or 32° Fahrenheit) and the standard pressure is 1 atm pressure.

pV = nRT

n = pV / RT

1 Pa = 1 Nm

R = 8.31432×103 N⋅m⋅kmol−1⋅K−1 = 8.31432 m3⋅Pa⋅K−1⋅mol−1

n (He) = 98 x 2.5 / (8.314 x (15+273)) = 0.10 mol


V (He)stp= nRT / p = (0.10 x 8.314 x 273) / 1 = 222 m3